OpenStudy (anonymous):

Word problem help, need to set up 2 separate linear equations............... A brick layer and an electrician together spend 90 hours working on a house. If the brick layer earns $12 per hour and the electrician earns $16 and the owner paid them a total of $1350, how many hours did each work?

OpenStudy (anonymous):

If x is the number of hours that the electrician works and y is the number of hours that the brick layer works, then the total number of hours worked is 90. x + y = 90. If x is the number of hours that the electrician works, then he will receive $16x. If y is the number of hours that the brick layer works, then he will receive $12y. The total amount of money is $1350 so... 16x+12y = 1350

OpenStudy (anonymous):

how did you get the numbers 16 and 12?

OpenStudy (anonymous):

The question said that the brick layer is paid $12 per hour, and the electrician $16. So, the total mount of money that the brick layer gets is 12 * the number of hours he worked, and for the electricial it is 16 * the number of hours worked.

OpenStudy (anonymous):

so my two equations would be x + y = 90 and 12x + 16y = 1350 ?

OpenStudy (anonymous):

Yup, exactly.